# EgoBlind: Towards Egocentric Visual Assistance for the Blind

### v2

@@ Abstract
We present EgoBlind, the first egocentric VideoQA dataset collected from blind individuals to evaluate the assistive capabilities of contemporary multimodal large language models (MLLMs). EgoBlind comprises 1,392 videos that record the daily lives of real blind users from a first-person perspective. It also features 5,311 questions directly posed or generated and verified by blind individuals to reflect their in-situation needs for visual assistance under various scenarios. We provide each question with an average of 3 reference answers to alleviate subjective evaluation. Using EgoBlind, we comprehensively evaluate 16 advanced MLLMs and find that all models struggle, with the best performers achieving accuracy near 60%, far behind human performance of 87.4%. To guide future advancements, we identify and summarize major limitations of existing MLLMs in egocentric visual assistance for the blind and explore heuristic solutions for improvement. With these efforts, we hope EgoBlind can serve as a valuable foundation for developing more effective AI assistants to enhance the independence of the blind individuals’ lives. Data and evaluation code are available at https://github.com/doc-doc/EgoBlind.
@@ 1 Introduction
The rapid advancement of multimodal large language models (MLLMs) [1, 2, 3, 4, 5, 6] has significantly improved the performance of visual question answering (VQA). However, existing VQA datasets primarily focus on third-person perspective [7, 8] or general-purpose image and video understanding [9, 10, 11], overlooking critical real-world applications such as visual assistance for the visually-impaired who constitute an overwhelming population of 2.2 billion people around the world [12]. The research is especially scarce in assisting the blind from a first-person perspective [13].
@@
In light of this, we construct EgoBlind – the first egocentric VideoQA dataset designed to benchmark and advance MLLMs towards live egocentric visual assistance for the blind. EgoBlind comprises 1,392 egocentric videos that capture the visual experiences of blind individuals and 5,311 questions that are directly posed or automatically generated and further verified by blind users to reflect their assistive needs in exploring their surroundings. To categorize these needs for better analysis, the questions are grouped into six key types: “Information Reading”, “Safety Warning”, “Navigation”,“Social Communication”, “Tool Use”, and “Other Resources”. We set the QA task as real-time (timestamp restricted) and open-form answer generation to better align with its live assistance nature. Multiple reference answers with well-aligned evaluation prompts are also provided for effective assessment.
Figure 1: Real example of EgoBlind about a blind user seeking egocentric visual assistance. The first time when she puts her hand on the knob of the microwave, she is curious about what the knob control. The second time after she tunes the knob, she is interested to know the position of the knob. The third time when she puts her hand on another knob, she is curious about what knob it is. Finally, she wants to know if it is correct to turn the knob clock-wise, to set the cooking time.
@@
By focusing on blind users’ needs for real-time egocentric visual assistance, we aim to explore several key challenges in video-language learning, including egocentric dynamic scene understanding with poor visual quality (e.g., unstable motion, object blur, and occlusions), in-situation user intention reasoning, assistive and reliable answer generation. Take Figure 1 as an example; engaging in the user’s activity from a first-person perspective is crucial to understanding and answering the questions. This specification requires capturing the gaze area and hand motion and reasoning the related visual content spatio-temporally. Finally, this should be achieved with poor quality visual inputs, as the regions of interest are often not in focus (e.g., the control panel in Figure 1).
@@
With EgoBlind, we benchmark 16 recent MLLMs, covering both the open-source models (e.g., InternVL2.5 [14], Qwen2.5-VL [4], LLaVA-OV [15] and VideoLLaMA3 [16]) and the closed-source ones (e.g., GPT-4o [5] and Gemini [6]). For open-source models, we specially take into account the models that 1) achieve the state-of-the-art (SOTA) results on common video QA benchmarks [17, 8, 18, 19], and benchmarks emphasizing egocentric understanding [11, 10, 20]. Additionally, we include three modes (Video-LLaVA [21] and LLaMA-VID [22], VILA1.5 [23]) that achieve SOTA on the image blind QA dataset VizWiz [24].
[…]
  • Models are poor in assisting with egocentric navigation, safety warning, and tool use (open-source ones), indicating limited knowledge in these fields.
@@
  • Models can generate objectively accurate answers relative to the visual contents, but the answers often fail to meet blind users’ subjective needs for assistance, indicating a weakness in reasoning about user intentions in specific situations.
  • Models are struggling in understanding the spatio-temporal contexts and localizing objects relative to the users’ real-time locations.
@@
To summarize our major contributions, 1) we tackle significant challenges to construct the first VideoQA dataset (EgoBlind) to benchmark and promote research toward egocentric visual assistance for the blind, and 2) we comprehensively analyze the behaviors of the leading MLLMs, revealing lots of limitations and possibilities for improvements.
@@ 2 Related Work
VQA for the Blind. Visual question answering (VQA) is the task of answering user questions about images and videos. As one of its most impactful applications, VQA has been proposed to help visually impaired individuals [26]. Yet, so far the vast majority of advancements have been made in general-purpose VQA [7, 27, 8, 17] where the questions are to test the models’ capability of general-purpose visual understanding, with also the images and videos captured from third-person perspective. The closest in aim is VizWiz [24], which collects data from real human-powered VQA systems for helping blind users (e.g., BeMyEye [^1] http://www.bemyeyes.org/). Despite the significance, VizWiz limits in static images and object-centric questions; it fails to cater to the real-time and broader assistive needs of blind users in exploring the dynamic surroundings. Recent ongoing work VIEW-QA [13] proposes to capture the daily challenges faced by visually impaired individuals via using 360-degree egocentric wearable cameras. However, its videos and questions are collected by training actors to simulate the experiences of blind individuals, limiting its ability to reflect the real blind people’s daily lives and their true needs for visual assistance, especially for those who have been blind for a long time.
@@
Egocentric VQA. Egocentric VQA has earned increasing interest for its significant application value towards embodied assistance [28, 29]. Early effort EgoVQA [30] is small scale and limits its questions in challenging action recognition from first-person perspective with the absence of the camera wearer in the footage. Subsequent advancement EgoTaskQA [9] uses machine-generated questions to evaluate models’ task understanding capabilities. AssistQ [31] is close to our aim for embodied assistance, but it limits to instructional videos for demonstration of tool use. Ever since the release of Ego4D [32], lots of real-world egocentric QA tasks are explored [33, 34, 35, 11, 36, 37, 10, 20]. However, most of they aim at general-purpose egocentric visual understanding, or focus on a specific aspect of assistance, e.g., QAEgo4D [35] for episodic memory, EgoTextVQA [38] for scene text understanding and EgoLifeQA [39] for long context life assistants. To our best knowledge, there is so far no existing egocentric VQA dataset specially collected from real blind individuals.
@@
MLLMs for VQA. MLLMs have significantly improved the landscape of VQA. By integrating visual information into powerful LLMs through lightweight connection modules (e.g., MLP [40]), MLLMs extend the capabilities of LLMs to converse with images and videos naturally like humans. This brings breakthrough over traditional VQA which answers questions by either multi-choice selection [41, 8] or close-vocabulary classification [27, 17]. While most of the MLLMs are developed for VQA of third-person perspective [29, 18, 42, 21, 43, 22, 44, 4, 16], recent advancements [33, 34, 45, 46, 14, 20] specially include ego videos for understanding. Nonetheless, all of they target a general-purpose egocentric visual understanding. In this paper, we will comprehensively examine their capacities towards egocentric visual assistance for the blind users.
@@ 3.1 Dataset Creation
Initially, we have tried to invite blind volunteers to record their daily lives from a first-person perspective and pose questions about their needs for daily visual assistance. However, we find that this approach presents significant challenges regarding data diversity (we can only get contact to a very limited number of blind people), quality (people often sparsely and coarsely ask what surrounds them to complete the data collection task), as well as ethical and privacy concerns. As an alternative, we collect videos from video media platforms and adopt the annotation pipeline as illustrated in Figure 2. We elaborate on the details below.
3.1.1 Video Collection
@@
We search and crawl 478 long-form egocentric videos (in batches) from video media platforms such as Bilibili and TikTok, ensuring that all videos are originated from visually impaired content creators. These videos, captured from the first-person perspective of blind individuals using GoPro or mobile phone, document their daily lives (such as in traveling, cooking, navigating, shopping, social gatherings, medical visiting.) and provide real-world data on how visually impaired individuals perform tasks and solve problems in dynamic environments. Since a single long video often contains a vast array of visual content, e.g., a summarized record of a visually impaired creator’s week, we manually parition the videos into distinct segments following the traces of splicing. During editing, we delete video moments with sharp scene transitions and big subtitles to enhance the focus of each clip. After editing, we obtain 1,329 video clips with an average duration of 40.0 seconds.
3.1.2 Question Collection
@@
We obtain the questions via three different approaches: 1) Manual Extraction: We manually extract the visually-assistive questions directly posed by the blind users in the videos. 2) AI-Generation: We prompt GPT-4o [5] to act as blind users and generate questions by engaging in their egocentric perspective. The generated questions are further verified and edited by both sighted and visually impaired people, ensuring their correctness and alignment with blind users’ true needs. 3) Blind User Annotation: For some videos, we introduce their major contents to the blind individuals and invite them to annotate assistive questions when confronted with the visual situations. The actual annotation is done together with the verification stage. Our collection details are as follows:
@@
Manual Extraction. We first manually watch the video clips and extract the vision-assistive questions from the camera wearers, the timestamps of the questions are also recorded. It is worth noting that these in-video questions are answered by the blind users’ sighted partners or others in the videos. Thus, to prevent answer leakage and ensure real-time QA, models can only access visual content up to the question timestamps to answer a specific question. We obtain a total of 541 questions during this process since qualified questions are sparse in the videos.
@@
AI-Generation. To enrich the annotations, we consider generating questions by prompting GPT-4o [5] to act as blind questioners, followed by rigorous verification with real blind individuals. Specifically, we first categorize the assistance scenarios into six types based on our observation of the video content and the extracted questions: Information Reading, Safety Warning, Navigation, Social Communication, Tool Use, and Other Resources. Typical question of each group is listed in Table 1. To ensure a diverse set of examples for each assistance type, the generation is conducted categorically with tailored prompt for each question type. In addition, we also prompt GPT to generate a reference answer for each question.
@@
Specifically, in our prompting for each category, the generated QA pairs are required to adhere to the following criteria to ensure that they are practical, context-aware, and closely aligned with the needs of the blind: 1) First-Person Perspective: All questions should be framed from the perspective of the visually impaired individuals (camera wearers). 2) Practicality: The questions must be practical and reflect the blind’s true needs in the given situation. 3) Video-Level Questions: Questions should reflect the temporal aspect of the video, requiring multiple frames to derive the answer and emphasizing dynamic characteristics. 4) Real-Time Context: Questions must pertain to the real-time context of the video and exclude any external information beyond the visible content up to the question time stamp. Other details are presented in Appendix A.1.
@@
3.1.3 Manual Verification
@@
Our verification is done in three stages involving both sighted and visually impaired people. The first three authors are engaged into all stages to control the quality. Concretely, in the first stage, the authors are responsible to check the QA quality every after generating a limited number of questions for one type of assistance. Related issues are recorded and reflected in the prompts for subsequent generations, thus to maximally ensure the generation quality. In such an alternate way, we generate 16,560 questions in batches.
@@
In the second stage, we invite 63 volunteers (sighted people) to thoroughly review the questions. In this process, we remove lots of questions that are 1) redundant in meaning; 2) not easy to answer (e.g., “How do I use my phone?”); and 3) meaningless (e.g., “How do I use my white cane?”). We also find and delete a moderate number of questions that are 1) not related to vision (e.g., “How heavy is this box?”) and 2) not from first-person perspective (e.g., “what is the man (camera wearer) doing?”) . Note that all ambiguous questions are further cross-validated by the three authors to either remove or edit them. What remained are 5,080 questions after this stage, in which more than 80% of data are edited by humans. Additionally, the 541 extracted questions also undergo checking for translation errors, resulting in a total of 5,621 valid questions after this stage.
@@
In the third stage, to reduce the annotation burden for the blind, we sample 30% (1.5K) questions and invite a total of 111 visually impaired volunteers to review them to ensure their relevance to the blind users’ true needs for visual assistance (details are presented in Appendix A.2). The result observations are then reflected throughout the whole dataset by the authors. Specifically, the blind volunteers are requested to score the confidence (from the lowest 0 to the highest 5) of a question that it may be asked under a specific visual scenario. Each question will be scored by at least 10 volunteers and maximally 33. Additionally, the bind volunteers are encouraged to pose new questions optionally. With the study result, we further remove the questions with an average confidence score of lower than 1 and add a limited number of new questions (about 300) annotated by the blind volunteers. Eventually, we obtain 5,311 valid questions. Their distributions are listed in Table 1.
3.1.4 Multiple Answer Annotation
@@
We invite 21 university students to watch the videos and answer the questions. To enhance the persuasiveness of the standard answers and reduce subjectivity, each question is assigned to three or four annotators for answering. Only one may see the generated answer (for generated questions only) and must verify and edit it as needed. The others are required to directly answer the question for assistance purpose. If there is no generated answer, e.g., for the manually extracted or annotated question, the annotators need to answer the question themselves. We resolve contradictory answers through consensus between the authors and the majority of annotators. Redundant answers are avoided by merging similar responses. Ultimately, each question has 1 to 4 valid answers (3 on average). Over 67% of the generated answers are edited by human annotators, and most unedited answers are deterministic, such as yes/no and numbers. Examples of answers are shown in Table 1.
3.2 Statistic Analysis
Table 2: Statistics of the EgoBlind datasets
Videos Questions
Train Test All Train Test All
673 656 1,329 2,746 2,565 5,311
Data Split. The data collection and annotation process takes about 10 months. We finalize 5,311 questions across 1,329 videos. For effective and efficient evaluation, we split half of the videos (656) along with their QA pairs (2,565) as test set, while leaving the remaining for instruction tuning towards model development. The split ensures that video clips from the same source video do not appear in both sets. Additionally, we employ Gemini 2.0 to probe if there is a significant discrepancy of model behaviors between the test set and the whole dataset. We find that the prediction accuracy differences are quite small, ranging from 0.2% to 0.5% across different question types. This suggests that the test set is sufficient to serve as a reliable evaluation ground for MLLMs. The detailed numbers are presented in Table 2.

### v3

@@ Abstract
We present EgoBlind, the first egocentric VideoQA dataset collected from blind individuals to evaluate the assistive capabilities of contemporary multimodal large language models (MLLMs). EgoBlind comprises 1,392 first-person videos from the daily lives of blind and visually impaired individuals. It also features 5,311 questions directly posed or verified by the blind to reflect their in-situation needs for visual assistance. Each question has an average of 3 manually annotated reference answers to reduce subjectiveness. Using EgoBlind, we comprehensively evaluate 16 advanced MLLMs and find that all models struggle. The best performers achieve an accuracy near 60%, which is far behind human performance of 87.4%. To guide future advancements, we identify and summarize major limitations of existing MLLMs in egocentric visual assistance for the blind and explore heuristic solutions for improvement. With these efforts, we hope that EgoBlind will serve as a foundation for developing effective AI assistants to enhance the independence of the blind and visually impaired. Data and code are available at https://github.com/doc-doc/EgoBlind.
@@ 1 Introduction
The rapid advancement of multimodal large language models (MLLMs) [1, 2, 3, 4, 5, 6] has significantly improved the performance of visual question answering (VQA). However, existing VQA datasets primarily focus on the third-person perspective [7, 8] or general-purpose image and video understanding [9, 10, 11, 12, 13, 14]. Applications such as visual assistance for the visually-impaired [15], who constitute an overwhelming 2.2 billion people around the world [16] have been received less attention. Research in assisting the blind from a first-person perspective is especially scarce [17].
@@
In light of this, we construct EgoBlind – the first egocentric VideoQA dataset designed to benchmark and advance MLLMs towards egocentric visual assistance for the blind. EgoBlind comprises 1,392 egocentric videos that capture the visual experiences of blind individuals and 5,311 questions that are directly posed or automatically generated and verified by blind users to reflect their assistive needs in exploring their surroundings. To categorize these needs for better analysis, the questions are grouped into six key types: “Information Reading”, “Safety Warning”, “Navigation”,“Social Communication”, “Tool Use”, and “Other Resources”. We set the QA task as online (timestamp restricted) and open-form answer generation to better align with its live assistance nature. Multiple reference answers with well-aligned evaluation prompts are also provided for effective assessment.
Figure 1: Example from EgoBlind about a blind user demonstrating egocentric visual assistance. As she places her hands on various microwave dials, she asks a series of questions about what the dial controls, its position and settings and how to adjust it.
@@
Our focus on blind users’ needs for egocentric visual assistance allows us to explore several key challenges in video-language learning, including egocentric dynamic scene understanding with poor visual quality (e.g., unstable motion, object blur, and occlusions), in-situation user intention reasoning, assistive and reliable answer generation. Take Figure 1 as an example; engaging in the user’s activity from a first-person perspective is crucial to understanding and answering the questions. This specification requires capturing the gaze area, hand motions and reasoning about related visual content spatio-temporally. Finally, this should be achieved with poor quality visual inputs, as the regions of interest may be out of focus or occluded (e.g., the control panel in Figure 1).
@@
With EgoBlind, we benchmark 16 recent MLLMs, covering the advanced open-source models (e.g., InternVL2.5 [18], Qwen2.5-VL [4]) and the closed-source ones (e.g., GPT-4o [5] and Gemini [6, 19]). For open-source models, we consider models that 1) achieve the state-of-the-art (SOTA) on common video QA benchmarks [20, 8, 21, 22], and egocentric understanding benchmarks [11, 10, 23]. Additionally, we include 3 models (Video-LLaVA [24] and LLaMA-VID [25], VILA1.5 [26]) that achieve SOTA on the blind image QA dataset VizWiz [15].
Our experimental results show that all models struggle on EgoBlind. The best performer (GPT-4o) achieves an accuracy of 59.3% and falls behind human performance of 87.4% by a whopping $\sim$28%. Interestingly, close-sourced models (e.g., Gemini 1.5 and 2.0), which often surpass open-source models in general-purpose VQA [8, 11, 22] perform worse than top-performing open-source models (e.g., InternVL2.5). Also, models superior on egocentric VQA are not necessarily the best-performing. Our further analyses and investigations lead to the following primary observations:
  • No single model wins across all assistance types, and most models are poor in egocentric navigation, safety warning, and communication, indicating research deficiency in these areas.
@@
  • Models may correctly answer questions about the visual scene, though the answers may fail to meet users’ needs for assistance, indicating a weakness in reasoning about user intentions.
  • Models struggle in reasoning the change of spatial orientation relative to the users from the sequence of egocentric visual inputs.
@@
  • Finetuning with EgoBlind training data effectively benefits model performances, though a large gap remains compared to human capabilities.
Our work tackles significant challenges to construct the first VideoQA dataset (EgoBlind) to benchmark and promote research toward egocentric visual assistance for the blind. We have comprehensively analyzed the behaviors of the leading MLLMs, revealing limitations and areas for improvements. With these efforts, we hope that EgoBlind will serve as a foundation for developing effective AI assistants to enhance the independence of the blind and visually impaired.
@@ 2 Related Work
VQA for the Blind. Visual question answering (VQA) is the task of answering user questions about images and videos [28, 29, 30]. One promising application area is to support the visually impaired. Yet, the majority of advancements have been made in general-purpose settings [7, 31, 8, 20] with images and videos captured from third-person perspectives. The closest in aim is VizWiz [15], which collects data from real human-powered VQA systems for helping blind users (e.g., BeMyEye [^1] http://www.bemyeyes.org/). VizWiz opens up the potential for visual assistance of the blind, but is limited in handling static images and object-centric questions; it fails to cater to the real-time and broader assistive needs of blind users in exploring the dynamic surroundings. Recent work VIEW-QA [17] captures the daily challenges faced by visually impaired individuals via using 360-degree egocentric wearable cameras. However, its videos and questions are collected from seeing actors who simulate the experiences of the blind, and thus may not reflect the daily lives and true needs of the blind for visual assistance, especially for those who have been blind for a long time.
@@
Egocentric VQA. Egocentric VQA has gained interest for its application value towards embodied assistance [32, 29]. Early effort EgoVQA [33] is small scale and limits its questions in challenging action recognition from first-person perspective with the absence of the camera wearer in the footage. Subsequent advancement EgoTaskQA [9] uses machine-generated questions to evaluate models’ task understanding capabilities. AssistQ [34] is close to our aim for embodied assistance, but it limits to instructional videos for demonstration of tool use. Ever since the release of Ego4D [35], lots of real-world egocentric QA tasks are explored [36, 37, 38, 11, 39, 40, 10, 23]. However, they mostly aim at general-purpose egocentric visual understanding, or focus on a specific aspect of assistance, e.g., QAEgo4D [38] for episodic memory, EgoTextVQA [41] for scene text understanding and EgoLifeQA [42] for long context life assistants. To our best knowledge, there is so far no existing egocentric VQA dataset specially collected from real blind individuals.
@@
MLLMs for VQA. By integrating visual information into powerful LLMs through lightweight connection modules (e.g., MLP [43]), MLLMs extend the capabilities of LLMs to converse with images and videos naturally like humans. MLLMs have significantly improved the landscape of VQA. This brings breakthrough over traditional VQA which answers questions by either multi-choice selection [44, 8] or close-vocabulary classification [31, 20]. While most MLLMs are developed for VQA from third-person views [29, 21, 45, 24, 46, 25, 47, 4, 48], recent advancements [36, 37, 49, 50, 18, 23] specifically add egocentric videos for understanding. Nonetheless, all of they target a general-purpose egocentric visual understanding. In this paper, we will comprehensively examine their capabilities towards egocentric visual assistance for the blind users.
@@ 3 EgoBlind Dataset
3.1 Video Collection
@@
We collect videos from video sharing platforms such as Bilibili and TikTok and design an annotation pipeline as illustrated in Figure 2. Specifically, we download 478 long-form egocentric videos (in batches) of blind and visually impaired content creators. These videos, captured using GoPro or mobile phones, document their daily lives while traveling, cooking, navigating through town, shopping, and in social gatherings. It provides near-real-world data on how visually impaired individuals perform daily tasks and solve problems in dynamic environments. Since a single long video often contains an array of visual content, e.g., a summarized record of a visually impaired creator’s week, we manually partition the videos into distinct segments based on the source edits. We exclude video moments with sharp scene transitions or strong staging and production indications (e.g., large subtitles, visual special effects, etc.) to focus primarily on life-logging. After partioning and filtering, we obtain 1,329 video clips with an average duration of 40 seconds.
3.2 Question Collection
@@
We obtain the questions via three different approaches: 1) Manual Extraction: We extract visually-assistive questions posed by the blind content creators in their videos. 2) AI-Generation: We prompt GPT-4o [5] to act as a blind user and generate questions by engaging in their egocentric perspective. The generated questions are further verified and edited by both seeing and blind people, ensuring their correctness and alignment with blind users’ true needs. 3) Blind User Annotation: For some videos, we describe the contents to blind annotators and ask them pose assistive questions as if they were in such a setting. The actual annotation is done together with the verification stage for AI generated questions. Our collection details are as follows:
@@
Manual Extraction We first manually watch the video clips and extract the vision-assistive questions from the camera wearers, the timestamps of the questions are also recorded. It is worth noting that these in-video questions are answered by the blind users’ sighted partners or others in the videos. Thus, to prevent answer leakage and ensure real-time QA, models can only access visual content up to the question timestamps to answer a specific question. We obtain a total of 541 questions during this process since qualified questions are sparse in the videos.
@@
AI-Generation. To enrich the annotations, we consider generating questions by prompting GPT-4o [5] to act as blind questioners, followed by rigorous verification with real blind individuals. We categorize the assistance scenarios into six groups based on our observation of video contents and the extracted questions: Information Reading, Safety Warning, Navigation, Social Communication, Tool Use, and Other Resources. Typical questions from each group is listed in Table 1. To ensure a diverse set of examples for each assistance type, the generation is conducted categorically with tailored prompts for each question type. In addition, we also prompt GPT to generate a reference answer for each question.
@@
In our prompts, the generated QA pairs are ensured to be context-aware and aligned with the needs of the blind: 1) Ego-centric: All questions should be framed from the perspective of the visually impaired individual (camera wearers). 2) Practical: The questions must reflect practical needs of the blind in the given situation. 3) Video-Level Questions: Questions should reflect temporal events in the video, emphasizing dynamic characteristics and requiring multiple frames to answer. 4) Online Context: Questions must pertain to the online context of the video and exclude any external information beyond the visible content up to the question time stamp. Other details are presented in the Appendix A.1.
@@
3.3 Manual Verification and Answer Annotation
@@
Our verification is done in three stages involving both seeing and visually impaired people. The first three authors are engaged in all stages for quality control. Concretely, in the first stage, the authors check the QA quality after generating a limited number of questions for each assistance type. Related issues are recorded and reflected in the updated prompts for subsequent generations to improve the generation quality. In such an alternate way, we generate 16,560 questions in batches.
@@
In the second stage, we invite 63 volunteers seeing individuals to thoroughly review the questions. In this process, we remove questions that are 1) redundant in meaning; 2) vague and difficult to answer directly (e.g., “How do I use my phone?”); and 3) meaningless (e.g., “How do I use my white cane?”). We also remove a moderate number of questions that are 1) not related to visual inputs (e.g., “How heavy is this box?”) and 2) not engaged in first-person perspective (e.g., “what is the man (camera wearer) doing?”). Note that all ambiguous questions are further cross-validated by the three authors to either remove or edit them. After this stage, 5,080 questions remained; more than 80% of them have been human-edited by this point. Additionally, the 541 extracted questions also undergo a check for language translation errors, resulting in a total of 5,621 valid questions after this stage.
@@
In the third stage, we sample 30% (1.5K) of the questions and invite 111 blind volunteers to review them for relevance to blind users’ needs for visual assistance. The sampling are visual situation based to cover all visual situations in EgoBlind (see details in Appendix A.2). Specifically, the blind volunteers are requested to score (from the lowest 0 to the highest 5) the confidence that a question would be asked under a specific visual scenario. Each question is scored by 10 to 33 volunteers. Additionally, the bind volunteers are encouraged to pose additional new questions. Afterwards, we remove the questions with an average confidence score lower than 1, and add approximately 300 new questions directly from the blind volunteers (Blind User Annotation). The result observations (what question to be deleted or added) are then reflected throughout the whole dataset by the authors. Finally, we obtain 5,311 valid questions. Their distributions are listed in Table 1.
@@
Multiple Answer Annotation. For each question, we annotate multiple ground-truth answers by inviting 21 university students to watch the related video. To enhance the persuasiveness of the standard answers and reduce subjectivity, each question is assigned to three or four annotators for answering. Only one may see the generated answer (for generated questions only) and must verify and edit it as needed. The others are required to directly answer the question for assistance purpose. If there is no generated answer, e.g., for the manually extracted or annotated question, the annotators need to answer the question themselves. We resolve contradictory answers through consensus between the authors and the majority of annotators. Redundant answers are avoided by merging similar responses. Ultimately, each question has 1 to 4 valid answers (3 on average). Over 67% of the generated answers are edited by human annotators, and most unedited answers are deterministic, such as yes/no and numbers. Examples of answers are shown in Table 1.
@@
3.4 Statistic Analysis
Table 2: Statistics of the EgoBlind datasets
Videos Questions
Train Test All Train Test All
673 656 1,329 2,746 2,565 5,311
Data Split. The data collection and annotation process takes about 10 months. We finalize 5,311 questions across 1,329 videos. For effective and efficient evaluation, we split half of the videos (656) along with their QA pairs (2,565) as a test set, while leaving the remaining for instruction tuning towards model development. The split ensures that video clips from the same source video do not appear in both sets (we also study a user-specific data-split strategy in Appendix B.3). Additionally, we use Gemini 2.0 to probe if there is a significant discrepancy of model behaviors between the test set and the whole dataset. We find that the prediction accuracy differences are quite small, ranging from 0.2% to 0.5% across different question types. This suggests that the test set is sufficient for reliable evaluation. The detailed numbers are presented in Table 2.